i have no idea how to start
X + 2y = -1
2x - y = 8
solve the system

Answer by checkley75(3666)

2X-Y=8
-Y=-2X+8
Y=2X-8 NOW REPLACE Y WITH (2X-8) IN THE OTHER EQUATION & SOLVE FOR X
X+2(2X-8)=-1
X+4X-16=-1
5X=-1+16
5X=15
X=15/5
X=3 ANSWER.
Y=2*3-8
Y=6-8
Y=-2 ANSWER.
PROOF
2*3-(-2)=8
6+2=8
8=8
-----------------------------------------------------------
OR YOU COULD USE THE ADDITION/SUBSTRATION METHOD:
MULTIPLY THE SECOND EQUATION BY 2 THUS:
4X-2Y=16 NOW ADD TO THE OTHER EQUATION:
X+2Y=-1
---------------------------------
5X=15
X=15/5
X=3 ANSWER.
3+2Y=-1
2Y=-1-3
2Y=-4
Y=-4/2
Y=-2 ANSWER.